# Adaptive Cascading Network for Continual Test-Time Adaptation

### v1

@@ Abstract.
We study the problem of continual test-time adaption where the goal is to adapt a source pre-trained model to a sequence of unlabelled target domains at test time. Existing methods on test-time training suffer from several limitations: (1) Mismatch between the feature extractor and classifier; (2) Interference between the main and self-supervised tasks; (3) Lack of the ability to quickly adapt to the current distribution. In light of these challenges, we propose a cascading paradigm that simultaneously updates the feature extractor and classifier at test time, mitigating the mismatch between them and enabling long-term model adaptation. The pre-training of our model is structured within a meta-learning framework, thereby minimizing the interference between the main and self-supervised tasks and encouraging fast adaptation in the presence of limited unlabelled data. Additionally, we introduce innovative evaluation metrics, average accuracy and forward transfer, to effectively measure the model’s adaptation capabilities in dynamic, real-world scenarios. Extensive experiments and ablation studies demonstrate the superiority of our approach in a range of tasks including image classification, text classification, and speech recognition.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
Intelligent systems operating in real-world settings frequently encounter non-stationary data distributions that evolve over time. For example, self-driving cars would face dynamically changing environments due to weather, lighting conditions, geographic locations, etc. This variability poses significant challenges to machine learning models as a highly accurate model on training data may fail catastrophically on shifted distributions (Goodfellow et al., 2015). These challenges necessitate model retraining to assimilate new distributions, also known as continual learning (McCloskey and Cohen, 1989). However, most existing continual learning approaches are designed primarily for fully labeled datasets, which is prohibitively expensive.
@@
Test-time adaptation (Sun et al., 2020; Wang et al., 2021; Prabhudesai et al., 2023) is a promising approach to adapt a pre-trained model to unlabelled target data without access to training data. The essence of these methods lies in the incorporation of a self-supervised learning (SSL) task, e.g., predicting image rotation, minimizing the entropy loss, etc. Recently, Wang et al. (Wang et al., 2022) has extended the problem setting to continual test-time adaptation, where there is a sequence of target distributions at test time, to better reflect the real-world scenario. However, recent work Liu et al. (Liu et al., 2021) shows that the unconstrained model update from the SSL task may interfere with the main task. This interference results in the accumulation of prediction errors and a gradual deviation from the model’s true predictive mechanism, preventing long-term model adaptation. Moreover, existing methods lack the ability of quickly adapting (e.g., adapting with few gradient steps) the model to the current distribution and demand statistically sufficient data (e.g. large batch size). This issue is crucial for continual and non-stationary settings when test data arrive in an online manner with small batches. In addition, test-time training methods (Sun et al., 2020; Liu et al., 2021; Bartler et al., 2021) typically employ a parallel paradigm (Fig. 1 (a)) with an auxiliary classifier for the SSL task, thus only enabling feature update while keeping the main classifier fixed at test time. Continually updating the model in such a manner would lead to a growing misalignment between the feature extractor and the main classifier, leading to deteriorated test accuracy.
@@ 2. Related Work
Meta-learning. Meta-learning (Schmidhuber, 1987) is a long-standing topic on learning models to generalize over a distribution of tasks. Model-Agnostic Meta-Learning (MAML) by Finn et al. (Finn et al., 2017) was proposed to adapt the model to new tasks within a few gradient steps. The key idea is to learn a good initialization from which the model is able to be quickly adapted to new tasks with few-shot examples. Several approaches have been proposed to use meta-learning to address distributional shifts. Li et al. (Li et al., 2017) suggested an episodic training paradigm to improve models’ generalization capability. Balaji et al. (Balaji et al., 2018) came up with MetaReg that meta-learned a regularization function that can generalize to new domains. Dou et al. (Dou et al., 2019) proposed to incorporate global and local constraints to learn semantic feature spaces in a modified MAML framework. MT3 (Bartler et al., 2021) was formulated to incorporate meta-learning into test-time adaptation by learning task-specific model parameters for different tasks. However, MT3 requires the meta-model to be accessed at test time which makes continual adaptation for a single model unsuitable.
@@ 4.2. Model Pre-training
Domain Randomization. First, we create simulated non-stationary target domains that the model would encounter during test time to support the meta-learning scheme. We employ domain randomization (Volpi et al., 2021) to generate domain augmentations from the single source $\mathcal{S}$ through a set of transformations $\Psi$. Given $\Psi$ and the source samples $\{(\bm{x}_{s},\bm{y}_{s})\}\sim\mathbb{P}_{\mathcal{S}}(X,Y)$ , we generate augmentations by sampling a specific transformation $\pi\sim\Psi$, and then applying it to the given data, obtaining $\left\{\left(\pi\left(\bm{x}_{s}\right),\bm{y}_{s}\right)\right\}\sim p(\mathcal{S}^{+})$. For each iteration, we sample a mini-batch $\tau\sim\mathbb{P}_{\mathcal{S}^{+}}(X,Y)\tau\sim\mathbb{P}_{\mathcal{S}^{+}}(X,Y)$ representing a random domain. We then divide $\tau$ into meta-train data $\mathcal{D}_{\tau}^{\text{trn}}$ and meta-validation data $\mathcal{D}_{\tau}^{\text{val}}$ for the meta-learning pre-training phase. Domain randomization varies across different modalities. The implementation details are provided in Sec. 6 and Appendix A.
@@ 6.1. Image Classification
Table 2. Results (%) on CIFAR-10-C on the gradually changing setup. Models are pre-trained on CIFAR-10 and continually adapted to a sequence of 135 gradually changing domains with a batch size of 32. Our method enables long-term model adaptation to target domains while others fail catastrophically.
@@ 6.4. Ablation Study
Furthermore, we apply meta-learning on TTT and achieve a significant improvement in performance across different batch sizes, as shown in Tab. 7. Similar to Eq. 3 and Eq. 4 described in Sec. 4.2, we adapt the inner and outer loop update to the TTT framework respectively as:
@@
where $\mathcal{L}_{\text{R}}$ is the image rotation prediction loss. Noticeably, it boosts TTT’s accuracy with the batch size of 128 by $4.11\%$. Another critical observation is that without meta-learning our cascading paradigm exhibits a poor performance for small batch sizes (16, 32 and 64) and has to rely on sufficient data statistics to yield a decent accuracy. Results demonstrate the effectiveness of meta-learning in mitigating the interference between the main and SSL tasks and facilitating fast adaptation with few unlabelled samples.
@@ 6.5. Uncertainty Quantification
In our approach, we use entropy as the self-supervised loss. Entropy can be used to quantify the uncertainty. To investigate the relationship between uncertainty and classification error, we analyze the evolution of entropy and error. Results are shown in Fig. 4. In Fig. 4 (left), we note that both entropy and error decrease when the model adapts to more samples of the target domain. In Fig. 4 (right), we can see that both entropy and error increase with the level of severity increasing. Entropy shows consistent trends with the classification error in both cases. The results demonstrate that entropy can reflect the model’s uncertainty with respect to target domains as well as measure the distributional distance between the source and target domain.

### v2

@@ Abstract.
We study the problem of continual test-time adaption where the goal is to adapt a source pre-trained model to a sequence of unlabelled target domains at test time. Existing methods on test-time training suffer from several limitations: (1) Mismatch between the feature extractor and classifier; (2) Interference between the main and self-supervised tasks; (3) Lack of the ability to quickly adapt to the current distribution. In light of these challenges, we propose a cascading paradigm that simultaneously updates the feature extractor and classifier at test time, mitigating the mismatch between them and enabling long-term model adaptation. The pre-training of our model is structured within a meta-learning framework, thereby minimizing the interference between the main and self-supervised tasks and encouraging fast adaptation in the presence of limited unlabelled data. Additionally, we introduce innovative evaluation metrics, average accuracy and forward transfer, to effectively measure the model’s adaptation capabilities in dynamic, real-world scenarios. Extensive experiments and ablation studies demonstrate the superiority of our approach in a range of tasks including image classification, text classification, and speech recognition. Our code is publicly available at https://github.com/Nyquixt/CascadeTTA.
@@
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679801isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
Intelligent systems operating in real-world settings frequently encounter non-stationary data distributions that evolve over time. For example, self-driving cars would face dynamically changing environments due to weather, lighting conditions, geographic locations, etc. This variability poses significant challenges to machine learning models as a highly accurate model on training data may fail catastrophically on shifted distributions (Goodfellow et al., 2015; qiao2021uncertainty; qiao2023topology; peng2022out; li2023data). These challenges necessitate model retraining to assimilate new distributions, also known as continual learning (McCloskey and Cohen, 1989). However, most existing continual learning approaches are designed primarily for fully labeled datasets, which is prohibitively expensive.
@@
Test-time adaptation (Sun et al., 2020; Wang et al., 2021; Prabhudesai et al., 2023; qiao2024ensemble) is a promising approach to adapt a pre-trained model to unlabelled target data without access to training data. The essence of these methods lies in the incorporation of a self-supervised learning (SSL) task, e.g., predicting image rotation, minimizing the entropy loss, etc. Recently, Wang et al. (Wang et al., 2022) has extended the problem setting to continual test-time adaptation, where there is a sequence of target distributions at test time, to better reflect the real-world scenario. However, recent work Liu et al. (Liu et al., 2021) shows that the unconstrained model update from the SSL task may interfere with the main task. This interference results in the accumulation of prediction errors and a gradual deviation from the model’s true predictive mechanism, preventing long-term model adaptation. Moreover, existing methods lack the ability of quickly adapting (e.g., adapting with few gradient steps) the model to the current distribution and demand statistically sufficient data (e.g. large batch size). This issue is crucial for continual and non-stationary settings when test data arrive in an online manner with small batches. In addition, test-time training methods (Sun et al., 2020; Liu et al., 2021; Bartler et al., 2021) typically employ a parallel paradigm (Fig. 1 (a)) with an auxiliary classifier for the SSL task, thus only enabling feature update while keeping the main classifier fixed at test time. Continually updating the model in such a manner would lead to a growing misalignment between the feature extractor and the main classifier, leading to deteriorated test accuracy.
@@ 2. Related Work
Meta-learning. Meta-learning (Schmidhuber, 1987) is a long-standing topic on learning models to generalize over a distribution of tasks. Model-Agnostic Meta-Learning (MAML) by Finn et al. (Finn et al., 2017) was proposed to adapt the model to new tasks within a few gradient steps. The key idea is to learn a good initialization from which the model is able to be quickly adapted to new tasks with few-shot examples. Several approaches (qiao2020learning; ma2021smil) have been proposed to use meta-learning to address distributional shifts. Li et al. (Li et al., 2017) suggested an episodic training paradigm to improve models’ generalization capability. Balaji et al. (Balaji et al., 2018) came up with MetaReg that meta-learned a regularization function that can generalize to new domains. Dou et al. (Dou et al., 2019) proposed to incorporate global and local constraints to learn semantic feature spaces in a modified MAML framework. MT3 (Bartler et al., 2021) was formulated to incorporate meta-learning into test-time adaptation by learning task-specific model parameters for different tasks. However, MT3 requires the meta-model to be accessed at test time which makes continual adaptation for a single model unsuitable.
@@ 4.2. Model Pre-training
Domain Randomization. First, we create simulated non-stationary target domains that the model would encounter during test time to support the meta-learning scheme. We employ domain randomization (Volpi et al., 2021; qiao2020learning) to generate domain augmentations from the single source $\mathcal{S}$ through a set of transformations $\Psi$. Given $\Psi$ and the source samples $\{(\bm{x}_{s},\bm{y}_{s})\}\sim\mathbb{P}_{\mathcal{S}}(X,Y)$ , we generate augmentations by sampling a specific transformation $\pi\sim\Psi$, and then applying it to the given data, obtaining $\left\{\left(\pi\left(\bm{x}_{s}\right),\bm{y}_{s}\right)\right\}\sim p(\mathcal{S}^{+})$. For each iteration, we sample a mini-batch $\tau\sim\mathbb{P}_{\mathcal{S}^{+}}(X,Y)\tau\sim\mathbb{P}_{\mathcal{S}^{+}}(X,Y)$ representing a random domain. We then divide $\tau$ into meta-train data $\mathcal{D}_{\tau}^{\text{trn}}$ and meta-validation data $\mathcal{D}_{\tau}^{\text{val}}$ for the meta-learning pre-training phase. Domain randomization varies across different modalities. The implementation details are provided in Sec. 6 and Appendix A.
@@ 6.1. Image Classification
Table 2. Results (%) on CIFAR-10-C on the gradually changing setup. Models are pre-trained on CIFAR-10 and continually adapted to a sequence of 135 gradually changing domains with a batch size of 32. Our method enables long-term adaptation to target domains while others fail catastrophically.
@@ 6.4. Ablation Study
Furthermore, we apply meta-learning on TTT and achieve a significant improvement in performance across different batch sizes, as shown in Tab. 7. Similar to Eq. 3 and Eq. 4 described in Sec. 4.2, we adapt the inner and outer loop update to TTT respectively as:
@@
where $\mathcal{L}_{\text{R}}$ is the image rotation prediction loss. Noticeably, it boosts TTT’s accuracy with the batch size of 128 by $4.11\%$. Another critical observation is that without meta-learning our cascading paradigm exhibits a poor performance for small batch sizes and has to rely on sufficient data statistics to yield a decent accuracy. Results demonstrate the effectiveness of meta-learning in adapting to few samples.
@@ 6.5. Uncertainty Quantification
In our approach, we use entropy as the self-supervised loss, which is used to quantify the uncertainty. To investigate the relationship between uncertainty and classification error, we analyze the evolution of entropy and error. Results are shown in Fig. 4. In Fig. 4 (left), we note that both entropy and error decrease when the model adapts to more samples of the target domain. In Fig. 4 (right), we can see that both entropy and error increase with the level of severity. Entropy shows consistent trends with the classification error in both cases. The results demonstrate that entropy can reflect the model’s uncertainty with respect to target domains as well as measure the distributional distance between the source and target domain.
@@ 7. Conclusion
Acknowledgements.
This work is supported by the National Science Foundation through the Faculty Early Career Development Program (NSF CAREER) Award NSF-IIS-2340074 and the Department of Defense under the Defense Established Program to Stimulate Competitive Research (DoD DEPSCoR) Award.